OpenStudy (anonymous):

Determine the intervals over which the function is increasing, decreasing, or constant. Find the location (x value) of any relative minimums or relative maximums. f(x)=-3x^3+x

OpenStudy (anonymous):

Please help!

OpenStudy (anonymous):

You can do it using algebra f(x)=-3x^3+x=x(1-3x^2)=x(1+sqrt3 x)(1-sqrt x) and checking this and interpreting it But it is easier if you just take the derivative and know that when f'(x)=0 the function has min/max

OpenStudy (anonymous):

The easiest way would be to graph the function and then look when its increasing, decreasing and constant

OpenStudy (anonymous):

You can do it using algebra f(x)=-3x^3+x=x(1-3x^2)=x(1+sqrt3 x)(1-sqrt x) and checking this and interpreting it But it is easier if you just take the derivative and know that when f'(x)=0 the function has min/max

OpenStudy (anonymous):

whatt are yall talking abiut. ?

OpenStudy (anonymous):

Thank you very much!